# Determination of nuclear quadrupole moments of 25Mg, 87Sr, and 135,137Ba via configuration-interaction plus coupled-cluster approach

### v1

@@ IV Conclusion
In this work, we calculated the magnetic dipole hyperfine-structure constants and electric-field gradients for the $3s3p~^{3}\!P_{1,2}$ states of 24Mg, the $5s4d~^{1}\!D_{2}$ and $5s5p~^{3}\!P_{1,2}$ states of 87Sr, and the $6s5d~^{3}\!D_{1,2,3}$, $6s5d~^{1}\!D_{2}$, and $6s6p~^{1}\!P_{1}$ states of 135,137Ba. We utilized five distinct computational methods, namely CI+MBPT, CI+LCCSD, CI+CCSD, CI+LCCSDs, and CI+CCSDs, to derive the final recommended values along with their corresponding uncertainties. Our final recommended values for the magnetic dipole hyperfine-structure constants of these states are fully consistent with the measured results considering uncertainties.
By incorporating the measured electric quadrupole hyperfine-structure constants and our calculated electric-field gradients of these states, we determined the nuclear quadrupole moments of 25Mg, 87Sr, 135Ba, and 137Ba to be 0.203(2) b, 0.336(4) b, 0.161(2) b, and 0.246(4) b, respectively. The nuclear quadrupole moment of 25Mg determined by us is highly consistent with the value obtained from the muonic X-ray experiment. Our calculated nuclear quadrupole moment of 87Sr is consistent not only with the previously extracted values from the $5p_{3/2}$ state of 87Sr+ but also with those from the $5s5p$ $^{3}\!P^{o}_{1}$ and $5s5p$ $^{3}\!P^{o}_{2}$ states of 87Sr. Moreover, our nuclear quadrupole moments of 135,137Ba are in agreement with the earlier results derived from the $6p_{3/2}$ state of 135,137Ba+. Nevertheless, there are approximately 10% and 4% discrepancies between our nuclear quadrupole moments of 87Sr and 137Ba and the currently adopted values Pyykkö (2018) obtained from the $4d_{5/2}$ state of 87Sr+ and the $5d_{3/2,5/2}$ state of 137Ba+. Since our nuclear quadrupole moment values were obtained through the statistical averaging of multiple states of neutral atoms, we are confident that our final recommended values are both accurate and reliable. Based on a recently high-precision theoretical work Chakraborty et al. (2025), we think that the previous experimental measurement results of the $4d_{5/2}$ state of 87Sr+ need to be further verified. Given that the contribution of triple excitations is not currently taken into account in our coupled-cluster calculations and the three-body interaction potential is also neglected, there is still room for improvement in our method. We also think that it is still meaningful to further verify the nuclear quadrupole moments of 135,137Ba through other theoretical calculations or new measurements.

### v2

@@ IV Conclusion
In this work, we calculated the magnetic dipole hyperfine-structure constants and electric-field gradients for the $3s3p~^{3}\!P_{1,2}$ states of 24Mg, the $5s4d~^{1}\!D_{2}$ and $5s5p~^{3}\!P_{1,2}$ states of 87Sr, and the $6s5d~^{3}\!D_{1,2,3}$, $6s5d~^{1}\!D_{2}$, and $6s6p~^{1}\!P_{1}$ states of 135,137Ba. We utilized five computational methods, namely CI+MBPT, CI+LCCSD, CI+CCSD, CI+LCCSDs, and CI+CCSDs, to derive the final recommended values along with their corresponding uncertainties. Our final recommended values for the magnetic dipole hyperfine-structure constants of these states are fully consistent with the measured results considering uncertainties.
By combining the measured electric-quadrupole hyperfine-structure constants with our calculated electric-field gradients for the relevant states, we obtain nuclear quadrupole moments of 0.203(2) b, 0.336(4) b, 0.161(2) b and 0.246(4) b for 25Mg, 87Sr, 135Ba, and 137Ba, respectively. Our value for 25Mg agrees with the muonic X-ray result considering the uncertainty. The 87Sr quadrupole moment is consistent with earlier extractions from the $5p_{3/2}$ state of 87Sr+ as well as from the the $5s5p$ $^{3}\!P^{o}_{1}$ and $5s5p$ $^{3}\!P^{o}_{2}$ states of 87Sr, and the nuclear quadrupole moments of 135,137Ba agree with those derived from the $6p_{3/2}$ state of 135,137Ba+. Nevertheless, there are approximately 10% and 4% discrepancies between our nuclear quadrupole moments of 87Sr and 137Ba and the currently adopted values Pyykkö (2018) obtained from the $4d_{5/2}$ state of 87Sr+ and the $5d_{3/2,5/2}$ state of 137Ba+. These discrepancies between neutral-atom- and ion-based extractions remain to be clarified by further experimental and theoretical works. The recent work indicates that the triple excitations contribute $1-2\%$ to the electric-quadrupole hyperfine-structure constants of the $d_{3/2}$ and $d_{5/2}$ states in Sr+ and Ba+ Chakraborty et al. (2026). To enhance further the accuracy of nuclear quadrupole moments derived with the RCI+CC calculations, the triple excitation contributions need to be included.